I do not understand how they got K*= 40

From 40^3 = 40K^2

• 9 = 300E1/3K1/3, 9. = 12,000, w = $10, r = $400, E*,K* = ? In gptimum (MRTSI = or comerc MPE ZE = 100 E ** MP =100 100 l
Answer #1

Answer : 40^3 = 40K^2

=> 40 * 40 * 40 = 40K^2

=> (40 * 40* 40) / 40 = K^2

=> 40 * 40 = K^2

=> 40^2 = K^2

=> K = √(40^2)

=> K = (40^2)^1/2

=> K = 40^(2 * 1/2)

=> K = 40

Thus here K = 40 comes from 40^3 = 40K.